OpenStudy (anonymous):

what is 9/2 + (-3/2) i dont know to to solve it

OpenStudy (anonymous):

Plus a negative is the same as minus! 9/2 - 3/2 = 6/2 since you already have common denominators, you can just do what the top says. The top is 9 - 3 which =6. Put that over the common denominator. Simple as that :)

OpenStudy (anonymous):

Or, if you like decimals .. 9/2 = 4.5 and 3/2 = 1.5 4.5 - 1.5 = 3 which equals 6/2 :)

OpenStudy (anonymous):

so the ans is 3??

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

thx